OpenStudy (anonymous):

solve for the complete square x^2+4x-8=0

jimthompson5910 (jim_thompson5910):

\[\large x^2+4x-8=0\] \[\large x^2+4x=8\] \[\large x^2+4x+4=8+4\] \[\large (x+2)^2=12\] \[\large x+2=\pm\sqrt{12}\] \[\large x+2=\pm2\sqrt{3}\] \[\large x=-2\pm2\sqrt{3}\]

OpenStudy (anonymous):

b/2^2 wasnt used..

jimthompson5910 (jim_thompson5910):

that's the point where I added 4 to both sides

jimthompson5910 (jim_thompson5910):

take half of 4 (the x coefficient) and square it to get 4, then add it to both sides

OpenStudy (anonymous):

okay i understand thanks

OpenStudy (whpalmer4):

We are trying to get left hand side to be in the form \[(x + a)^2 = c\] or \[x^2 + 2ax + a^2 = c\] \[2a=4, a=2\] \[(x+2)^2 = x^2 + 4x + 4\] So we need to add 12 from both sides to make the left side equal to (x+2)^2, giving us \[(x+2)^2 = 12\] which means \[(x+2) = \pm\sqrt(12)\] and \[x=\pm2\sqrt(3)-2\]